# references.bib

### v1

@@ 5.4 Research Questions
Answer to RQ2: Empirical analysis of GitHub resources allowed us to build a meaningful taxonomy. The expert interviews were indeed crucial to complementing and re-structuring our initial taxonomy.
Surveys corroborated the results of repository mining and interviews and also provided additional (initial) insights into the severity of the faults and the level of effort involved in resolving them.
The taxonomy’s structure, organised around categories corresponding to the fault’s location, and sub-categories corresponding to its cause is relevant to both fault detection and localisation, and will easily adapt to potential evolutions of hybrid architectures in the future.
6 Discussion
@@
6.1 Hybrid Quantum-Classical faults vs. Faults in Deep Learning Systems
@@
Hybrid Quantum-Classical systems bear some similarities to deep learning systems [rieser_tensor_2023][beer_training_2020]; thus, it is natural to compare our taxonomy with similar work in machine learning [humbatova_taxonomy_2019]. While several categories may seem similar at first, noticeably Optimisation, API, and GPU, the parallel is superficial: the root causes in those categories remain different in Hybrid Quantum-Classical architectures. Other categories are unique to the Hybrid Quantum-Classical architectures. More specifically, for quantum machine learning (QML), although the optimisation process of the parametric gates is similar to training in deep learning, the encoding of the dataset into qubits is specific to QML and thus, the faults introduced in this architecture cannot be directly related to faults in deep learning.
6.2 Computer Science vs. Physics Expertise
Quantum computing is a multi-disciplinary field. While computer scientists and physicists use the same architectures, they approach them differently and experience different problems. This is noticeable in the interviews. Computer scientists typically struggle with conceptualisation, designing the Ansatz and mapping a problem to a Hamiltonian. Physicists struggle with the optimisation phase.
6.3 Manual Debugging Methods
@@
Two approaches were mentioned to investigate and resolve the identified faults: one focusing on the optimisation process, and another on the problem conceptualisation. In the first method, the developer detects non-convergence, and first focuses on debugging the implementation and behaviour of the gradients. They would manually debug each gradient at every iteration to check if the results are correctly passed and to detect where the error may originate from. The second approach is more problem-specific and was mentioned by several experts with a Physics background. In case of a wrong behaviour of the system, they would first focus on their problem definition and manually ensure that the Hamiltonian properly represents the problem. The next step would be to manually ensure that the Hamiltonian is correctly mapped into qubits. Mapping a Hamiltonian into qubits is not trivial, and finding an efficient mapping is an open question [tilly_variational_2022]. They would then focus on the structure of the Ansatz, and ensure no electronic symmetry is broken (see Section 5.2). Only then, if no bugs are found in the previous steps, they would proceed to similar steps as the first approach. In this interview, the expert gives a detailed explanation of such a procedure. Many of these steps could be automated in the future.
6.4 Common Problems
@@
The semi-structured interviews led us to discuss experts’ problems beyond our taxonomy. A topic mentioned frequently was the lack of accurate and up-to-date documentation. Quantum platforms tend to mostly document their basic functionalities, leaving advanced features undocumented. Moreover, experts miss a flexible Hybrid Quantum-Classical architecture to experiment on. Tensorflow Quantum [broughton_tensorflow_2021] was introduced with this purpose, but stopped being maintained. Several interviewees mentioned having to implement their connections between machine learning frameworks and quantum simulators, which often introduce compatibility errors and performance issues. With the growing interest in hybrid architecture, researchers and developers would benefit from a dedicated platform that would strongly integrate all components of such architectures and ensure robust and reliable implementations, particularly where quantum and classical parts interact. Although common platforms such as Qiskit and Pennylane allow some support for common NISQ algorithms, such as VQA, experts lack more flexible architecture and templates for various components - typically Ansatz and Optimisation, with clear guidelines for the potential applications and customisations.
7 Threats to validity
@@
Hybrid Quantum-Classical architectures are relatively new, and their implementations undergo constant revisions. It is challenging to capture structural faults not solely caused by the evolving implementations. It is also challenging to know to what extent our taxonomy can be generalised to future architectures. While we expect that noise-related problems may be mitigated, other faults in this taxonomy inherent to the hybrid classical-quantum architecture should remain relevant.
8 Conclusions
@@
We plan to incorporate our dataset into a toolbox for generating faulty implementations of Hybrid Quantum-Classical architectures. This will serve as a means for evaluating future techniques for testing, quality assurance, and repair of such architectures, which is a direction of our ongoing research. We would like to further investigate other views that can be used to classify our dataset of real faults.

### v2

@@ 5.4. Research Questions
Answer to RQ2: Our taxonomy was developed iteratively through a combination of empirical mining and expert validation. The initial structure emerged from a detailed analysis of GitHub issues, where faults were manually labeled based on their symptoms and technical origin. Expert interviews then played a crucial role in refining the categories, reorganising them to better reflect the practical understanding of fault causes.
For instance, an issue that would initially have been labelled as a “Structural” fault happening at the Ansatz level was reclassified during interviews as a conceptual error in Ansatz design, drawing attention to deeper modelling challenges that are specific to Hybrid Quantum-Classical architectures. This shift highlighted the importance of differentiating between surface-level symptoms and root causes—an insight made possible by expert feedback. Moreover, the interviews revealed disciplinary nuances, such as computer scientists struggling more on conceptual faults and physicists on optimisation challenges, which helped balance the taxonomy to better capture the interdisciplinary nature of the field.
Importantly, the interviews also helped counterbalance the overrepresentation of computer science perspectives typically observed in online resources such as GitHub. Several interviewees, particularly from physics and mathematics backgrounds, explicitly mentioned not publishing or publicly reporting encountered faults—a bias that would otherwise limit the generalisability of the dataset. These qualitative insights ensured a more comprehensive and representative taxonomy.
The final taxonomy structure organises faults by their location (e.g., classical optimisation, quantum circuit, API layer) and further subcategorises them by cause (e.g., memory issue, circuit decomposition, suboptimal API integration). This approach ensures the taxonomy is both empirically grounded and practically relevant, supporting fault detection, localisation, and future adaptability to evolving Hybrid architectures.
6. Discussion
@@
6.1. Hybrid Quantum-Classical faults vs. Faults in Deep Learning Systems
@@
Hybrid Quantum-Classical systems bear some similarities to deep learning systems (Rieser et al., 2023)(Beer et al., 2020); thus, it is natural to compare our taxonomy with similar work in machine learning (Humbatova et al., 2019). While several categories may seem similar at first, noticeably Optimisation, API, and GPU, the parallel is superficial: the root causes in those categories remain different in Hybrid Quantum-Classical architectures. Other categories are unique to the Hybrid Quantum-Classical architectures. More specifically, for quantum machine learning (QML), although the optimisation process of the parametric gates is similar to training in deep learning, the encoding of the dataset into qubits is specific to QML and thus, the faults introduced in this architecture cannot be directly related to faults in deep learning.
6.2. Computer Science vs. Physics Expertise
Quantum computing is inherently a multi-disciplinary field. While computer scientists and physicists often use the same Hybrid Quantum-Classical architectures, their approaches and the types of problems they encounter can differ significantly. This difference is clearly reflected in our interviews. Computer scientists tend to struggle more with conceptualisation-related faults, such as designing the Ansatz or effectively mapping a problem to a Hamiltonian, where the physical properties and noise characteristics of quantum hardware play a critical role. Physicists, on the other hand, often face more challenges during the optimisation phase, which requires abstract reasoning about algorithms and software design.
These disciplinary differences also influence how faults are documented and shared within the community. Our interviews revealed that computer scientists are generally more likely to publish faults and open issues on public platforms like GitHub, which is reflected in the dominance of computer science perspectives in data mined from such sources. Conversely, experts from physics, chemistry, and mathematics backgrounds often do not publish faults or open issues publicly, either due to the sensitive nature of experimental setups or different cultural norms in their fields. This discrepancy was crucial to uncover through expert interviews, as it highlighted the limitations of relying solely on public datasets for taxonomy development and motivated our inclusion of a conceptualisation category based on the unique insights from these interviews. Overall, this underlines the importance of combining empirical mining with expert consultation to capture a fuller, more balanced picture of the fault landscape in Hybrid Quantum-Classical architectures.
6.3. Manual Debugging Methods
@@
Two approaches were mentioned to investigate and resolve the identified faults: one focusing on the optimisation process, and another on the problem conceptualisation. In the first method, the developer detects non-convergence, and first focuses on debugging the implementation and behaviour of the gradients. They would manually debug each gradient at every iteration to check if the results are correctly passed and to detect where the error may originate from. The second approach is more problem-specific and was mentioned by several experts with a Physics background. In case of a wrong behaviour of the system, they would first focus on their problem definition and manually ensure that the Hamiltonian properly represents the problem. The next step would be to manually ensure that the Hamiltonian is correctly mapped into qubits. Mapping a Hamiltonian into qubits is not trivial, and finding an efficient mapping is an open question (Tilly et al., 2022). They would then focus on the structure of the Ansatz, and ensure no electronic symmetry is broken (see Section 5.2). Only then, if no bugs are found in the previous steps, they would proceed to similar steps as the first approach. In Interview 5, the expert gives a detailed explanation of such a procedure. Many of these steps could be automated in the future.
6.4. Common Problems
@@
The semi-structured interviews led us to discuss experts’ problems beyond our taxonomy. A topic mentioned frequently was the lack of accurate and up-to-date documentation. Quantum platforms tend to mostly document their basic functionalities, leaving advanced features undocumented. Moreover, experts miss a flexible Hybrid Quantum-Classical architecture to experiment on. Tensorflow Quantum (Broughton et al., 2021) was introduced with this purpose, but stopped being maintained. Several interviewees mentioned having to implement their connections between machine learning frameworks and quantum simulators, which often introduce compatibility errors and performance issues. With the growing interest in Hybrid architecture, researchers and developers would benefit from a dedicated platform that would strongly integrate all components of such architectures and ensure robust and reliable implementations, particularly where quantum and classical parts interact. Although common platforms such as Qiskit and Pennylane allow some support for common NISQ algorithms, such as VQA, experts lack more flexible architecture and templates for various components - typically Ansatz and Optimisation, with clear guidelines for the potential applications and customisations.
6.5. Practical Applications and Guidelines for Use
Hybrid Quantum-Classical architectures are an emerging area of research with increasing practical significance. However, prior studies have yet to empirically capture the specific nature and limitations of faults in such architectures. Our work addresses this gap by introducing the first structured taxonomy of real-world faults in Hybrid Quantum-Classical architectures, grounded in actual implementations and validated through expert feedback. This constitutes a novel theoretical contribution to the field of quantum software engineering, offering a foundational understanding of fault landscapes in Hybrid architectures.
Beyond its theoretical value, our taxonomy is supported by a curated dataset of 133 empirically observed faults, enabling several practical applications:
  • Grounding testing and debugging tools in real-world hybrid faults, thus aligning tool development with the actual challenges faced by developers.
  • Enabling fault injection and simulation using empirically derived fault patterns to benchmark the effectiveness of quality assurance techniques.
  • Supporting education and documentation by illustrating recurring pitfalls in hybrid program implementation, thereby aiding onboarding and training.
  • Guiding future empirical studies through a reusable and extensible classification framework, facilitating comparative analysis across platforms and toolchains.
We plan to integrate this taxonomy and dataset into a toolbox for generating faulty implementations of Hybrid architectures. This resource will support the empirical evaluation of testing, debugging, and automated repair techniques. We are conducting a comprehensive analysis of our findings, derived from both the dataset and the interviews, to develop testing oracles capable of identifying incorrect behaviours that do not result in crashes for the different fault types by classifying their symptoms. These oracles constitute a critical step toward advancing the understanding of hybrid faults and provide a foundation for future research in Quantum Software Testing, both at the theoretical and empirical levels. By making both the taxonomy and dataset publicly available, we aim to promote reproducibility and accelerate the advancement of robust, testable Hybrid Quantum-Classical architectures.
Furthermore, this field is inherently interdisciplinary, drawing on deep expertise from both quantum physics and software engineering. As such, our taxonomy and dataset can serve as a bridge between communities: helping researchers with physics backgrounds gain insight into practical software engineering challenges, and conversely, assisting software engineers in understanding the fundamental constraints and behaviours of quantum systems when embedded into classical computations.
The expert interviews conducted as part of this study also provide valuable qualitative data, offering diverse perspectives on how Hybrid Quantum-Classical architectures are developed and maintained in practice. These interviews can serve as a resource for researchers and practitioners seeking a deeper understanding of the ecosystem, tooling gaps, and fault patterns encountered by different categories of users.
Practical Guidelines
To support both practitioners and researchers working with Hybrid Quantum–Classical architectures, we propose four main use cases for our taxonomy and dataset. These use cases illustrate how structured knowledge of causes (faults) and their effects (symptoms and failures) can drive improved system design, testing, and debugging practices.
  1. (1)
    Categorising Faults in Hybrid Architectures and Analysing Cause (Fault) — Effect (Symptom/Failure) Relationships: The taxonomy organises common failure causes across the life cycle of hybrid architectures, from parametrisation and conceptualisation to API, optimisation, and circuit implementation. For example, at the parametrisation stage, developers often introduce faults such as wrong input types for tensors or incorrect initial points in parametric initialisation; in the conceptualisation stage, faults caused by suboptimal Ansatz design or broken electronic symmetries can arise. By linking these causes to their typical effects (symptoms and failures), researchers can study relationships between specific mistakes and the effects they trigger. This knowledge can inform the design of better development practices, guide education efforts, and support future cause-prediction or fault-prediction models.
  2. (2)
    Developing Specialised Oracles to Trigger Effects Caused by Faults: The dataset supports the design of test oracles aimed at detecting causes that lead to silent but incorrect behaviours — that is, faults that do not cause program crashes. For instance, a fault in Ansatz design can produce a barren plateau that silently degrades optimisation performance, while high noise in measurement processes can cause biased outcomes without explicit error messages, and violations of physical constraints, such as particle number or spin, may go unnoticed. By implementing oracles that monitor gradient magnitudes, symmetry preservation, and statistical properties of outputs, and integrating them into continuous testing pipelines, silent failures can be detected early. These oracles improve reliability and provide actionable feedback for hybrid developers.
  3. (3)
    Test Case Generation Techniques: Insights from the taxonomy guide the development of input generation methods and test suite optimisation. Strategies include:
    • Parametric Variation: systematically vary initial tensor parameters, circuit depths, or optimizer hyperparameters to expose faults in parametrisation and optimisation phases.
    • Circuit Structure Manipulation: explore alternative Ansatz, qubit mappings, or distributed circuit decompositions to uncover implementation and parallelisation faults, including inconsistencies when recombining subcircuits.
    • Constraint Violation Probes: generate states that intentionally challenge physical constraints (e.g., particle number, spin, symmetry) to test the sensitivity and effectiveness of physics-consistency oracles.
    • Noise Stress Testing: inject controlled stochastic gate errors or simulate decoherence to quantify robustness and identify brittle components.
    • Hybrid Workflow Coverage: design sequences that combine classical preprocessing, quantum execution, and classical postprocessing in varied configurations to expose integration or optimizer-induced failures.
    Linking test generation strategies to cause categories enables practitioners to focus on the most fault-prone aspects of hybrid architectures and produce efficient, targeted, and reproducible test suites that go beyond naïve random sampling.
  4. (4)
    Automatic Fault Localisation and Repair: The structured mapping from faults to effects supports automated debugging tools. Observed failure patterns can be traced back to likely root causes. For example, a fault in parameter-shift rule implementation during optimisation often manifests as low-fidelity outputs, while memory-handling causes — such as improper GPU allocation — may cause performance degradation. Combined with specialised oracles and targeted test cases, this information supports semi-automatic or fully automated repair strategies, such as suggesting alternative Ansatz, modifying circuit decomposition strategies, or adjusting optimizer settings. This approach can shorten debugging cycles, improve reproducibility, and enhance software reliability in large-scale hybrid architectures.
Potential Application Areas
The proposed taxonomy and dataset are broadly applicable across key domains that rely on Hybrid Quantum–Classical architectures. These include: (i) quantum chemistry, where variational algorithms such as VQE are used to approximate molecular energies; (ii) drug discovery and repurposing, where hybrid workflows optimise quantum–classical pipelines for molecular similarity and binding prediction; (iii) optimisation and operations research, encompassing quantum approximate optimisation algorithms (QAOA) and hybrid solvers used for scheduling, routing, and resource allocation. In these contexts, systematic understanding of faults and their effects can improve robustness, reproducibility, and the interpretability of hybrid computational outcomes.
7. Threats to validity
@@
Hybrid Quantum-Classical architectures are relatively new, and their implementations undergo constant revisions. It is challenging to capture structural faults not solely caused by the evolving implementations. It is also challenging to know to what extent our taxonomy can be generalised to future architectures. While we expect that noise-related problems may be mitigated, other faults in this taxonomy inherent to the Hybrid Quantum-Classical architecture should remain relevant. Recent algorithmic paradigms such as the realisation of a quantum neural network using repeat-until-success circuits in a superconducting quantum processor (Moreira et al., 2023), exemplify the pace at which the field evolves. While these approaches introduce new design patterns, we believe that several fault types we documented—such as parameterization faults, quantum circuit design errors, and API misuse—are likely to remain relevant for such emerging models. Furthermore, similar fault types already appear in our dataset, particularly in contexts related to quantum machine learning. This suggests that certain structural weaknesses in Hybrid Quantum-Classical software development persist across algorithmic generations. Even though specific faults from these new paradigms were not directly observed during our study period, our taxonomy provides a principled framework that can help identify, classify, and reason about faults in these evolving algorithmic models.
8. Conclusions
@@
We plan to incorporate our dataset into a toolbox for generating faulty implementations of Hybrid Quantum-Classical architectures. This will serve as a means for evaluating future techniques for testing, quality assurance, and repair of such architectures, which is a direction of our ongoing research. We would like to further investigate other views that can be used to classify our dataset of real faults. As future work, we aim to build on these findings—both from the fault dataset and the interviews—to develop testing oracles capable of detecting incorrect behaviours that do not result in crashes. Such oracles, directly informed by this taxonomy, will represent an important step toward more reliable and robust Hybrid Quantum–Classical architectures.